OpenStudy (anonymous):

2. What is the geometric mean of 4 and 12? Use radicals when neede

ganeshie8 (ganeshie8):

multiply the numbers and take "square root"

OpenStudy (anonymous):

okay I got 48, how do I take the square root? Sorry, i'm pretty clueless with this.

563blackghost (563blackghost):

You would simply just find the square root of 48...so do you need this in radical form or decimal ?

OpenStudy (anonymous):

Radical form

563blackghost (563blackghost):

So to get it in radical form we would find two factors of 48 in which one is a perfect square....so in this case 16 and 3.... \(\Huge{\sqrt{16 \times 3}}\) What number multiples itself to equal 16?

OpenStudy (anonymous):

8

563blackghost (563blackghost):

Not quite. \(\Huge{8 \times 8 = 64 \neq 16}\)

OpenStudy (anonymous):

4x4=16 sorry I was thinking addition lol

563blackghost (563blackghost):

Correct ^^ Now that we know that the square root of 16 is 4 we would take the 4 to the outside of the root and leave the 3 there...so the radical form is... \(\Huge{4\sqrt{3}\)

563blackghost (563blackghost):

\(\Huge{4\sqrt{3}}\)*

OpenStudy (anonymous):

Okay, thank you so much for your help!

563blackghost (563blackghost):

np ^^